# Differentially private projection-depth-based medians

### v2

@@ Appendix B Proofs from Section 3
Let $U_{(j),u}=F_{\nu_{u}}((X^{\top}u)_{(j)})$. Now, recall that for any set of $n$ independent standard uniform random variables $V_{1},\ldots,V_{n}$, for any $j\in[n-1]$, it holds that $V_{(j+1)}-V_{(j)}\stackrel{{\scriptstyle\text{d}}}{{=}}V_{(1)}$ and $V_{(1)}\sim Beta(1,n)$ [see, e.g., 20]. These two facts, combined inequality (C.5), results in
@@ Appendix E Results concerning Condition 2
$\zeta_{n}(t)=e^{-cn(1-2\alpha)t^{2}/\lVert\Sigma^{-1/2}\rVert^{2}+d}.$
@@
We now focus on $I$, and apply an $\epsilon$-net argument. Recall that the set $\mathcal{N}_{\epsilon}$ is an $\epsilon$-net on $\mathbb{S}^{d-1}$ if for all $u\in\mathbb{S}^{d-1}$, there exists $v\in\mathcal{N}_{\epsilon}$ such that $\lVert u-v\rVert\leq\epsilon$. Letting $\mathcal{N}_{\epsilon}$ be an arbitrary $\epsilon$-net on $\mathbb{S}^{d-1}$, it is straightforward (see, e.g., [26, Exercise 4.4.4]) to show that for any $c>0$, $y\in\mathbb{R}^{d}$ and $20<\epsilon<1/$ 2,
@@
Furthermore, there exists $\mathcal{N}_{1/4}$ such that $|\mathcal{N}_{1/4}|\leq 9^{d}$ [26, see Corollary 4.2.13]. Applying these facts, along with Gaussian concentration of measure yields that there exists a universal constant $c>0$ such that

### v3

@@ Appendix B Proofs from Section 3
Let $U_{(j),u}=F_{\nu_{u}}((X^{\top}u)_{(j)})$. Now, recall that for any set of $n$ independent standard uniform random variables $V_{1},\ldots,V_{n}$, for any $j\in[n-1]$, it holds that $V_{(j+1)}-V_{(j)}\stackrel{{\scriptstyle\text{d}}}{{=}}V_{(1)}$ and $V_{(1)}\sim Beta(1,n)$ [see, e.g., 23]. These two facts, combined inequality (C.5), results in
@@ Appendix E Results concerning Condition 2
$\zeta_{n,\mu}(t)=\zeta_{n,\sigma}(t)=e^{-cn(1-2\alpha)t^{2}/\lVert\Sigma^{-1/2}\rVert^{2}+d}.$
@@
We now focus on $I$, and apply an $\epsilon$-net argument. Recall that the set $\mathcal{N}_{\epsilon}$ is an $\epsilon$-net on $\mathbb{S}^{d-1}$ if for all $u\in\mathbb{S}^{d-1}$, there exists $v\in\mathcal{N}_{\epsilon}$ such that $\lVert u-v\rVert\leq\epsilon$. Letting $\mathcal{N}_{\epsilon}$ be an arbitrary $\epsilon$-net on $\mathbb{S}^{d-1}$, it is straightforward (see, e.g., [30, Exercise 4.4.4]) to show that for any $c>0$, $y\in\mathbb{R}^{d}$ and $20<\epsilon<1/$ 2,
@@
Furthermore, there exists $\mathcal{N}_{1/4}$ such that $|\mathcal{N}_{1/4}|\leq 9^{d}$ [30, see Corollary 4.2.13]. Applying these facts, along with Gaussian concentration of measure yields that there exists a universal constant $c>0$ such that